Assume that we are auditing the records of Forde Corporation. A physical inventory has been taken by the company under our observation. However, the valuation extensions have not been completed. The records of the company show the following account data. The gross margin last period was 35% of net sales; we anticipate that it will be 25% for the year under audit.

Sales, gross $1,512,000 Beginning inventory $480,000
Sales returns (returned to inventory) 24,000 Freight-in 33,600
Purchases, gross 744,000 Purchase returns and allowances 9,600

Estimate the cost of ending inventory using the gross profit method. $Answer
Answer #1

>> Net Sales = $ 1512000 - $ 24000 = $ 1488000.

>> Gross profit = $ 1488000 * 25 % = $ 372000.

>> Net Purchases = $ 744000 - $ 9600 = $ 734400.

>> Closing Inventory = Opening inventory + Net Purchases + Freight in + Gross profit - Net Sales.

>> Closing Inventory = 480000 + 734400 + 33600 + 372000 - 1488000 = $ 132000.